40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 22 15:46:58 2014, Kurosawa, Summary, IOO, Seven transfer functions MC_OLTF.pdf
    Reply  Mon Dec 22 21:53:08 2014, rana, koji, Update, IOO, Seven transfer functions MC_OLG.pdf
       Reply  Tue Dec 23 01:55:35 2014, rana, koji, Update, IOO, Seven transfer functions MC_FSS_TF.pdf
          Reply  Tue Dec 23 20:50:39 2014, rana, koji, Update, IOO, Seven transfer functions MC_OLTF.pdfMC_OLTF2.pdfmatlab.zip
             Reply  Sun Jan 4 22:08:46 2015, rana, Update, IOO, MC loop characterizations: PZT/EOM crossover MC_OLGs.pdfMC_xover.pdf
             Reply  Mon Jan 5 23:36:47 2015, ericq, Update, IOO, AO cable reconnected 
Message ID: 10841     Entry time: Tue Dec 23 20:50:39 2014     In reply to: 10833     Reply to this: 10851   10855
Author: rana, koji 
Type: Update 
Category: IOO 
Subject: Seven transfer functions 

Today we decided to continue to modify the TTFSS board.

The modified schematic can be found here: https://dcc.ligo.org/D1400426-v1 as part of the 40m electronics DCC Tree.

What we did

1) Modify input elliptic filter (L1, C3, C4, C5) to give zero and pole at 30 kHz and 300 kHz, respectively. L1 was replaced with a 1 kOhm resistor.  C3 was replaced with 5600 pF. C4 and C5 were removed. So the expected locations of the zero and pole were at 28.4 kHz and 256 kHz, respectively. This lead filter replaces the Pomona box, and does so without causing the terrible resonance around 1 MHz.

2) Removed the notch filters for the PC and fast path. This was done by removing L2, L3, and C52.

At this point we tested the MC locking and measured the transfer function. We successfully turned up the UGF to 170kHz and two super-boosts on.

3) Now a peak at 1.7MHz was visible and probably causing noise. We decided to revert L2 and adjusted C50 to tune the notch filter in the PC path to suppress this possible PC resonance. Again the TF was measured. We confirmed that the peak at 1.7MHz is at -7dB and not causing an oscillation. The suppression of the peak is limited by the Q of the notch. Since its in a weird feedback loop, we're not sure how to make it deeper at the moment.

4) The connection from the MC board output now goes in through the switchable Test1 input, rather than the fixed 'IN1'. The high frequency gain of this input is now ~4x higher than it was. I'm not sure that the AD829 in the MC board can drive such a small load (125 Ohms + the ~20 Ohms ON resistance of the MAX333A) very well, so perhaps we ought to up the output resistor to ~100-200 Ohms?


Also, we modified the MC Servo board: mainly changed the corner frequencies of the Super Boost stages and some random cleanup and photo taking. I lost the connecting cable from the CM to the AO input (unlabeled).

  1.  The first two Super Boost stages were changed from 20k:1k to 10k:500 to give us back some phase margin and keep the same low freq gain. I don't really know what the gain requirement is for this servo here at the 40m. The poles and zeros were chosen for iLIGO so as to have the frequency noise be 10x less than the SRD at 7 kHz.
  2. The third Super Boost (which we never used) was changed from 10k:500 to ~3k:150 (?) just in case we want a little more low freq gain.
  3. There was some purple vestigial wiring on the back side of the board with a flying resistor; I think this was a way to put a DC offset in to the output of the board, but its not needed anymore so I removed it.

 

Attachment 1: MC_OLTF.pdf  859 kB  | Hide | Hide all | Show all
MC_OLTF.pdf
Attachment 2: MC_OLTF2.pdf  726 kB  | Show | Hide all | Show all
Attachment 3: matlab.zip  1.287 MB
ELOG V3.1.3-